a solution containing 121.8g of Zn(NO2)2 per liter has a density of 1.107 gmL^-1. calculate its molal concentration.
1
Expert's answer
2020-09-02T10:11:45-0400

Molal concentration of solution is 0.875 mol/kg

Cm =moles of solute/kg of solvent

V(solute)=121.8 g/1.107g/mL-1=116.35mL

V(solvent)= 1000mL-116.35mL=883.65 mL=0.88L=0.88 kg

n=m/M=121.8g/157.4g/mol=0.77 mol

Cm=0.77 mol/0.88 kg=0.875 mol/kg